OpenStudy (anonymous):

carl has 135 in the bank and plans to save 5 each week jean has 90 in the bank and plans to save 10 a week. how many weeks will it be before jean has at least as much in the bank as carl equation please?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

do the math

OpenStudy (anonymous):

but i have to show the work? there isnt just a simple equation?

OpenStudy (anonymous):

dude its not hard to do the math

OpenStudy (anonymous):

i didnt say it was dude

OpenStudy (anonymous):

135 + 5w = 90 + 10w 5w - 10w = 90 - 135 -5w = -45 w = 9 They will be equal at 9 weeks

OpenStudy (anonymous):

any questions ?